B&B has a new baby powder ready to market. If the firm goes directly to the market with the product, there is only a 65 percent chance of success. However, the firm can conduct customer segment research, which will take a year and cost $1.25 million. By going through research, B&B will be able to better target potential customers and will increase the probability of success to 80 percent. If successful, the baby powder will bring a present value profit (at time of initial selling) of $19.5 million. If unsuccessful, the present value payoff is $6.5 million. The appropriate discount rate is 13 percent.

  

Calculate the NPV for the firm if it conducts customer segment research and if it goes to market immediately. (Do not round intermediate calculations and enter your answers in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89.)
Answer #1

NPV of customer segment research:

NPV = -1,250,000 + (0.8*19,500,000 + 0.2*6,500,000) / (1 + 0.13)

NPV = -1,250,000 + (15,600,000 + 1,300,000] / 1.13

NPV = -1,250,000 + 14,955,752.21

NPV = $13,705,752.21

NPV if goes to market immediately:

NPV = 0.65*19,500,000 + 0.35*6,500,000

NPV = 12,675,000 + 2,275,000

NPV = $14,950,000.00